OpenStudy (anonymous):

The indefinite intergral of (x-1/3x)^2 dx equals

OpenStudy (anonymous):

\[\int\limits_{?}^{?}(x-1/3x)^2 dx\]

OpenStudy (amistre64):

if 3x is a demoniator; you might wanna start out by making the innards into a single term

OpenStudy (amistre64):

3x^2 -1 -------- ; and square it 3x 9x^4 -6x^2 +1 -------------- ; split it apart 9x^2 9x^4 6x^2 1 ----- (-) ----- (+) ----- ; now simplfy and integrate perhpas? 9x^2 9x^2 9x^2

OpenStudy (anonymous):

I got 1/3x^3-2/3x-1/(9x) But the choice given are: a. 1/3(x-1/(3x))^2 b.x^2-2/3+1/(9x^2) c.x^3/3-2x-1/(9x) d. x^3/2-2/3 x-9/x e.none of these can u help me double check?

OpenStudy (amistre64):

that is correct: \[\frac{1}{3}x^3-\frac{2}{3}x-\frac{1}{9x} +C\] \[\frac{3x^4-6x^2+C}{9x}\] unless there is some algebra that I cant see, the answer would appear to be e"none of the above".

OpenStudy (anonymous):

Okay thanks alot!!